OpenStudy (anonymous):

What is the value of the expression? 8! / 6!x2! A. 1 B. 28 C. 56 D. 3360

OpenStudy (anonymous):

@marissalovescats

OpenStudy (tkhunny):

General Rule: Never use "x" to mean multiplication. Very confusing. Have you considred writing out a few of the factors of "8!"?

OpenStudy (anonymous):

8! is 8*7*6*5*4*3*2*1

OpenStudy (tkhunny):

Well, the first 3 would have been sufficient. \(\dfrac{8!}{6!} = \dfrac{8\cdot 7\cdot 6!}{6!}\) Anything strike you about that?